# Complete complex Finsler metrics and uniform equivalence of the Kobayashi metric

### v3

@@ Abstract
In this paper, first of all, according to Lu’s and Zhang’s works about the curvature of the Bergman metric on a bounded domain and the properties of the squeezing functions, we obtain that Bergman curvature of the Bergman metric on a bounded strictly pseudoconvex domain with $C^{2}$-boundary or bounded convex domain is bounded. Secondly, by the property of curvature symmetry on a Kähler manifold, we have the property: if holomorphic sectional curvature of a Kähler manifold is bounded, we can deduce that its sectional curvature is bounded. After that, applying to the Schwarz lemma from a complete Kähler manifold into a complex Finsler manifold, we get that a bounded strictly pseudoconvex domain with $C^{2}$-boundary or bounded convex domain admit complete strongly pseudoconvex complex Finsler metric such that their holomorphic sectional curvature are bounded from above by a negative constant. Finally, by the Schwarz lemma from a complete Kähler manifold into a complex Finsler manifold, we prove the uniform equivalences of the Kobayashi metric and Carathéodory metric on a bounded strongly convex domain with smooth boundary.
Keywords: Complex Finsler metrics; Bounded domains; Squeezing fucntions; Holomorphic sectional curvature.
@@ 1 Introduction
Bounded domains are elementary objects of study in complex analysis. From the Riemann mapping theorem, we know that simply connected open subset in $\mathbb{C}$ which is not the whole complex plane is biholomorphic to the unit disk. In several complex variables, the situation is more complicated and mysterious. To study complex and geometric structures of bounded domains, one may consider that holomorphic mappings from bounded domains to some standard domains such as balls and vice versa. Deng et al. [7, 8] introduced the notion of squeezing function (see Definition 3.1 and 3.2) to study geometric and analytic properties of bounded domains.
@@
In 2012, Deng et al. [7] proved the squeezing function on any bounded domain is continuous. Soon after, they study boundary behaviors of the squeezing functions on some bounded domains. And they proved that the squeezing function of any strongly pseudoconvex domain tends to $1$ near the boundary. What’s more, for a bounded strongly pseudoconvex domain with $C^{2}$-boundary in $\mathbb{C}^{n}$, it is HHR. In 2016, Kim and Zhang [11] proved that the bounded convex domains admit the uniform squeezing property. Furthermore, all bounded convex domains in $\mathbb{C}^{n}(n\geq 1)$ are HHR. By the above results, we can know that the squeezing functions of bounded pseudoconvex domains with $C^{2}$-boundary or bounded convex domains in $\mathbb{C}^{n}$ are bounded from below by a positive constant.
@@
The terminology, HHR or USq, has been introduced in order to the study of completeness and equivalence of the classical invariant metrics, including the Teichmüller metric, the Bergman metric, the complete Kähler-Einstein metric of negative scalar curvature, the Carathéodory metric and the Kobayashi metric. The completeness of these metrics on various type of complex manifolds and the boundness of the holomorphic sectional curvature of the Bergman metric has been an interesting topic ever since 1926 (when the Carathéodory metric was first introduced [3]). It is known that the Kobayashi metric is the maximum metric among metric satisfying the decreasing property, however, the Carathédory metric is the minimum. It is known that on a bounded, smooth , strictly, pseudoconvex domain in $\mathbb{C}^{n}$, all classical invariant metrics are uniformly equivalent to each other (see, for example, [2, 15, 16, 28, 19, 8, 30] and the references therein).
@@
It has been well-known that for any bounded domain in $\mathbb{C}^{n}$, the holomorphic sectional curvature of the Bergman metric is always less than $2$ (see [12]). In 2015, Lu [17] obtained some lower bounds of the holomorphic sectional curvature and Ricci curvature of the Bergman metric on a bounded domain of $\mathbb{C}^{n}$. The methods rely on constructing proper minimal functions. We note that the lower bounds Lu obtained tends to $\infty$- ∞ at the point $z$ goes to the boundary $\partial D$. Based on Lu’s work, Zhang [31] studied the scalar curvature of the Bergman metric on a bounded domain of $\mathbb{C}^{n}$. They gave the lower and upper bounds estimates for the Bergman curvatures in terms of the squeezing function originally introduced by Deng et al. [7]. Combining the above resluts with Lu’s and Zhang’s work, we have the following theorem.
@@
(cf. Theorem 3.2) Suppose that $\mathcal{D}$ is a bounded pseudoconvex domain with $C^{2}$-boundary or bounded convex domain in $\mathbb{C}^{n}$. For each point $z\in\mathcal{D}$ and $v\in T^{1,0}_{z}\mathcal{D}\backslash\{0\}\cong\mathbb{C}^{n}\backslash\{0\}$, then
$C_{1}\leq Sec_{\mathcal{D}}(z,v)\leq C_{2},$
$C_{3}\leq Ric_{\mathcal{D}}(z,v)\leq C_{4},$
$C_{5}\leq Scal_{\mathcal{D}}(z,v)\leq C_{6},$
where $C_{1},C_{2},C_{3},C_{4},C_{5},C_{6}$ are only dependent in complex dimension $n$ of the complex manifold $\mathcal{D}$, $C_{1},C_{3},C_{5}$ are negative constants, $C_{2},C_{4},C_{6}$ are positive constants.
According to S. S. Chern, Finsler geometry is just Riemannian geometry without quadratic restrictions. In real Finsler geometry, there are many examples of non-quadratic Finsler metrics with various kind of properties such as curvature properties, completeness. In complex Finsler geometry, there are numerous strongly pseudoconvex complex Finsler metrics in the strict sense of Abate and Patrizio. In invariant metrics, the Kobayashi metric and the Carathéodry metric are complex Finsler metrics. Due to Lempert’s result (see Theorem 5.1 and Remark 5.1), the Kobayashi metric and the Carathéodory metric in bounded strongly convex with smooth boundary are complex Finsler metrics in sense of Abate and Patrizio [1]. But they are not explicit. Zhong [33] found there are non-Hermitian quadratic strongly pseudoconvex complex Finsler metrics on a domain of $\mathbb{C}^{n}$ in the sense of smooth category. It is very natural to ask the question: are these metrics complete? The following theorem answer this question.
Theorem 1.2.
(cf. Theorem 6.4) For a suitable choice of positive constant $C$, suppose that $\mathcal{D}$ is a bounded pseudoconvex domain with $C^{2}$-boundary or bounded convex domain in $\mathbb{C}^{n}$. Suppose that $\mathcal{D}$ admits a strongly pseudoconvex complex Finsler metric $G:T^{1,0}\mathcal{D}\rightarrow[0,+\infty)$ such that its holomorphic sectional curvature is bounded from above by a negative constant $-K$ and $g_{B}$ is the Bergman metric on $\mathcal{D}$. Then $H=CG+g_{B}$ and the Bergman metric $g_{B}$ are equivalent. What’s more, $CG+g_{B}$ is complete.
@@
[28, p. 665, Question 1] Does a (complete) Kobayashi-hyperbolic manifold $M$ admit a (complete) Hermitian metric of strongly negative holomorphic sectional curvature, in sense that there exists a negative constant $c$ such that the holomorphic sectional curvature $H$ satisfies $H<c$
@@
Recently, Nie and Zhong [21] proved that a bounded domain in $\mathbb{C}^{n}$ admits a non-Hermitian quadratic strongly pseudoconvex complex Finsler metric such that its holomorphic sectional curvature is bounded from above by a negative constant. Actually, these complex Finsler metrics are strongly convex. Combining Theorem 1.2 and the above result, we obtain the following theorem. Our result partly answer Question 1.1.
@@
(cf. Corollary 6.1) Suppose that $\mathcal{D}$ is a bounded pseudoconvex domain with $C^{2}$-boundary or bounded convex domain in $\mathbb{C}^{n}$. Then $\mathcal{D}$ admits a complete strongly pseudoconvex complex Finsler metric $H:T^{1,0}\mathcal{D}\rightarrow[0,+\infty)$ such that its holomorphic sectional curvature is bounded from above by a negative constant. What’s more, $\mathcal{D}$ is a complete Kobayashi-hyperbolic.
@@
[10, p. 112, Remark (2)] Let $(M,h)$ be a simply-connected complete Kähler manifold satisfying $-A\leq\text{sectional curvature}\leq-B$ for two positive constants $A$ and $B$. Then, the Kobayashi metric $\mathfrak{K}$ satisfies
@@
In [29], Wu and Yau confirmed this conjecture. In fact they proved a stronger result, as they relax the sectional curvature to the holomorphic sectional curvature, and remove the assumption of simply-connectedness.
@@
[29, p. 104, Theorem 2] Let $(M,h)$ be a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $-A\leq K_{h}\leq-B$ for some positive constants $A$ and $B$. Then the Kobayashi metric $\mathfrak{K}$ satisfies
@@
As point out in [10], it is well known that the left inequality in the conjecture follows from the Ahlfors-Schwarz lemma and the hypothesis of sectional curvature bounded above by a negative constant. It is very natural to ask the question: Does the right inequality in the conjecture follow from the Schwarz lemma and the hypothesis of holomorphic sectional curvature? Applying to the Schwarz lemma from a complete Kähler into a complex Finsler manifold [20, 21, 22, 18], we get the following theorem.
@@
(cf. Corollary 7.1) Suppose that $\mathcal{D}$ is a bounded strongly convex domain with smooth boundary in $\mathbb{C}^{n}$. And let $(\mathcal{D},h)$ be a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $-A\leq K_{h}\leq-B$ for some positive constants $A$ and $B$. Then Kobayashi metric and Carathéodory metric satisfy
@@
Here $C>0$ is only depending only $A,B$, and independent of complex dimension $n$ of complex manifold $\mathcal{D}$.
@@
(cf. Theorem 7.2) For a suitable choice of positive constant $C$, suppose that $(M,h)$ is a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $-A\leq K_{h}\leq B$ for some positive constants $A$ and $B$. If the Carathéodory metric $\mathfrak{C}^{2}$ on complex manifold $M$ is a smooth strongly pseudoconvex complex Finsler mtric. Then $H=C\mathfrak{C}^{2}+h$ and the Kähler metric $h$ are equivalent. What’s more, $M$ is a complete Kobayashi-hyperbolic manifold.
@@ 2 Terminology and preliminaries
In this section, we recall the following some basic definitions and facts on a complex Finsler manifold. For more details, we refer to [1, 13].
@@
is positive definte on $\widetilde{M}$.
Any $C^{\infty}$ Hermitian metric on a complex manifold $M$ is naturally a strongly pseudoconvex complex Finsler metric. Conversely, if a complex Finsler metric $G$ on a complex manifold $M$ is $C^{\infty}$ over the whole holomorphic tangent bundle $T^{1,0}M$, then it is necessary a $C^{\infty}$ Hermitian metric. That is, for any $(z;v)\in T^{1,0}M$
$G(z;v)=g_{\alpha\overline{\beta}}(z)v^{\alpha}\overline{v}^{\beta}$
@@
for a $C^{\infty}$ Hermitian tensor $g_{\alpha\overline{\beta}}$ on $M$. For this reason, in general the non-trivial (non-Hermitian quadratic) examples of complex Finsler metrics are only required to be smooth over the slit holomorphic tangent bundle $\widetilde{M}$. A manifold $M$ endowed with a strongly pseudoconvex complex Finsler metric $G$ is called a complex Finsler manifold.
@@
Let $(M,G)$ be a complex Finsler manifold $M$, and take $v\in\widetilde{M}$. Then the holomorphic sectional curvature $K_{G}(v)$ of $G$ along a non zero tangent vector $v$ is given by
@@
Abate and Patrizio found an phenomenon that the holomorphic sectional curvature sectional curvature of a complex Finsler metric $G$ is the supremum of the Gaussian curvature of the induced metric through a family of holomorphic maps.
@@
where the supremum is taken with respect to the family of all holomorphic maps $\varphi:\mathbb{D}\rightarrow M$ with $\varphi(0)=z$ and $\varphi^{\prime}(0)=\lambda v$ for some $\0\lambda\in\mathbb{C}\backslash\{0\}$; $K(\varphi^{*}G)(0)$ is the Gaussian curvature of $(\mathbb{D},\varphi^{*}G)$ at the point 00.
The Bergman pseudometric $g_{B}$ on a complex manifold $M$ of complex dimension $n$ is a Kähler pseudometric with local potential given by the coefficients of the Bergman kernel $K(x,x)$.
Let $f$ be a $L^{2}$-holomorphic $n$-form on $M$. In terms of local coordinates $\left(z_{1},\ldots,z_{n}\right)$ on a coordinate chart $U$, let $e_{K_{M}}=dz^{1}\wedge\ldots\wedge dz^{n}$ be a local basis of the canonical line bundle $K_{M}$ on $U$. We can write $f$ as $f_{U}e_{K_{M}}$ on $U$. Let $f_{i}$ be an orthonormal basis of $L^{2}$-sections in $H_{(2)}^{0}\left(M,K_{M}\right)$. Note that from conformality, the choice is independent of the metric on $M$. The Bergman kernel is given by $K(x,x)=\sum_{i}f_{i}\wedge\overline{f_{i}}$. Let $K_{U}(x,x)=\sum_{i}f_{U,i}\overline{f_{U,i}}$ be the coefficient of $K(x,x)$ in terms of the local coordinates. The Bergman metric is given by the Kähler form

### v4

@@ Abstract
In this paper, first of all, according to Lu’s and Zhang’s works about the curvature of the Bergman metric on a bounded domain and the properties of the squeezing functions, we observe that Bergman curvatures of the Bergman metric on a bounded strictly pseudoconvex domain with $C^{2}$-boundary or bounded convex domain are bounded. Applying to the Schwarz lemma from a complete Kähler manifold into a complex Finsler manifold, we get that a bounded strictly pseudoconvex domain with $C^{2}$-boundary or bounded convex domain admits complete strongly pseudoconvex complex Finsler metrics such that their holomorphic sectional curvature is bounded from above by a negative constant. Finally, by the Schwarz lemma from a complete Kähler manifold into a complex Finsler manifold, we prove the uniform equivalences of the Kobayashi metric and Carathéodory metric on a bounded strongly convex domain with smooth boundary.
Keywords: Complex Finsler metrics; Bounded domains; Squeezing functions; Holomorphic sectional curvature.
@@ 1 Introduction
Bounded domains are elementary objects of study in complex analysis. From the Riemann mapping theorem, we know that simply connected open subset in $\mathbb{C}$ which is not the whole complex plane is biholomorphic to the unit disk. In several complex variables, the situation is more complicated and mysterious. To study complex and geometric structures of bounded domains, one may consider that holomorphic mappings from bounded domains to some standard domains such as balls and vice versa. Deng et al. [7, 8] introduced the notion of squeezing function (see Definition 2.6 and 2.7) to study geometric and analytic properties of bounded domains.
@@
In 2012, Deng et al. [7] proved the squeezing function on any bounded domain is continuous. Soon after, they study boundary behaviors of the squeezing functions on some bounded domains. And they proved that the squeezing function of any strictly pseudoconvex domain tends to $1$ near the boundary. What’s more, for a bounded strictly pseudoconvex domain with $C^{2}$-boundary in $\mathbb{C}^{n}$, it is HHR. In 2016, Kim and Zhang [11] proved that the bounded convex domains admit the uniform squeezing property. Furthermore, all bounded convex domains in $\mathbb{C}^{n}(n\geq 1)$ are HHR. By the above results, we can know that the squeezing functions of bounded strictly pseudoconvex domains with $C^{2}$-boundary or bounded convex domains in $\mathbb{C}^{n}$ are bounded from below by a positive constant.
@@
The terminology, HHR or USq, has been introduced in order to the study of completeness and equivalence of the classical invariant metrics, including the Teichmüller metric, the Bergman metric, the complete Kähler-Einstein metric of negative scalar curvature, the Carathéodory metric and the Kobayashi metric. The completeness of these metrics on various types of complex manifolds and the boundness of the holomorphic sectional curvature of the Bergman metric has been an interesting topic ever since 1926 (when the Carathéodory metric was first introduced [3]). It is known that the Kobayashi metric is the maximum metric among metric satisfying the decreasing property, however, the Carathéodory metric is the minimum. It is known that on a bounded, smooth, strictly, pseudoconvex domain in $\mathbb{C}^{n}$, all classical invariant metrics are uniformly equivalent to each other (see, for example, [2, 15, 16, 30, 19, 8, 32] and the references therein).
@@
It has been well-known that for any bounded domain in $\mathbb{C}^{n}$, the holomorphic sectional curvature of the Bergman metric is always less than $2$ (see [12]). In 2015, Lu [17] obtained some lower bounds of the holomorphic sectional curvature and Ricci curvature of the Bergman metric on a bounded domain of $\mathbb{C}^{n}$. The methods rely on constructing proper minimal functions. We note that the lower bounds Lu obtained tends to -$\infty$ at the point $z$ goes to the boundary $\partial D$. Based on Lu’s work, Zhang [33] studied the scalar curvature of the Bergman metric on a bounded domain of $\mathbb{C}^{n}$. They gave the lower and upper bounds estimates for the Bergman curvatures in terms of the squeezing function originally introduced by Deng et al. [7]. Combining the above results with Lu’s and Zhang’s work, the holomorphic sectional curvature of the Bergman metric on a bounded strictly pseudoconvex domain with $C^{2}$-boundary or bounded convex domain is bounded.
In complex Finsler geometry, there are numerous strongly pseudoconvex complex Finsler metrics in the sense of Abate and Patrizio. In invariant metrics, the Kobayashi metric and the Carathéodory metric are complex Finsler metrics. Due to Lempert’s result (see Theorem 2.1 and Remark 2.1), the Kobayashi metric and the Carathéodory metric on bounded strongly convex with smooth boundary are complex Finsler metrics in sense of Abate and Patrizio [1]. But they are not explicit. Zhong [35] found there are non-Hermitian quadratic strongly pseudoconvex complex Finsler metrics on a domain of $\mathbb{C}^{n}$ in the sense of smooth category. It is very natural to ask the question: are these metrics complete ? The following theorem answers this question.
@@
(cf. Theorem 3.4) For a suitable choice of positive constant $C$, suppose that $\mathcal{D}$ is a bounded strictly pseudoconvex domain with $C^{2}$-boundary or bounded convex domain in $\mathbb{C}^{n}$. Suppose that $\mathcal{D}$ admits a strongly pseudoconvex complex Finsler metric $G:T^{1,0}\mathcal{D}\rightarrow[0,+\infty)$ such that its holomorphic sectional curvature is bounded from above by a negative constant K-K and $g_{B}$ is the Bergman metric on $\mathcal{D}$. Then $H=CG+g_{B}$ and the Bergman metric $g_{B}$ are equivalent. What’s more, $H=CG+g_{B}$ is complete.
@@
[30, p. 665, Question 1] Does a (complete) Kobayashi-hyperbolic manifold $M$ admit a (complete) Hermitian metric of strongly negative holomorphic sectional curvature, in sense that there exists a negative constant $c$ such that the holomorphic sectional curvature $H$ satisfies $H<c$ ?
@@
Recently, Nie and Zhong [21] proved that a bounded domain in $\mathbb{C}^{n}$ admits a non-Hermitian quadratic strongly pseudoconvex complex Finsler metric such that its holomorphic sectional curvature is bounded from above by a negative constant. Actually, these complex Finsler metrics are strongly convex. Combining Theorem 1.1 and the above result, we obtain the following theorem. Our result partly answers Question 1.1.
@@
(cf. Corollary 3.1) Suppose that $\mathcal{D}$ is a bounded strictly pseudoconvex domain with $C^{2}$-boundary or bounded convex domain in $\mathbb{C}^{n}$. Then $\mathcal{D}$ admits a complete strongly pseudoconvex complex Finsler metric $H:T^{1,0}\mathcal{D}\rightarrow[0,+\infty)$ such that its holomorphic sectional curvature is bounded from above by a negative constant. What’s more, $\mathcal{D}$ is a complete Kobayashi-hyperbolic.
@@
[10, p. 112, Remark (2)] Let $(M,h)$ be a simply-connected complete Kähler manifold satisfying $A\leq\text{sectional curvature}\leq-B$ for two positive constants $A$ and $B$. Then, the Kobayashi metric $\mathfrak{K}$ satisfies
@@
In [31], Wu and Yau confirmed this conjecture in the following theorem.
@@
[31, p. 104, Theorem 2] Let $(M,h)$ be a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $A\leq K_{h}\leq-B$ for some positive constants $A$ and $B$. Then the Kobayashi metric $\mathfrak{K}$ satisfies
@@
As point out in [10], it is well known that the left inequality in the conjecture follows from the Ahlfors-Schwarz lemma and the hypothesis of sectional curvature bounded above by a negative constant. It is very natural to ask the question: Does the right inequality in the conjecture follow from the Schwarz lemma and the hypothesis of holomorphic sectional curvature ? Applying to the Schwarz lemma from a complete Kähler into a complex Finsler manifold [21] (For more details about Schwarz lemma on complex Finsler manifolds, we refer to [20, 22, 18, 27, 23]), we get the following theorem.
@@
(cf. Corollary 4.1) Suppose that $\mathcal{D}$ is a bounded strongly convex domain with smooth boundary in $\mathbb{C}^{n}$. And let $(\mathcal{D},h)$ be a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $A\leq K_{h}\leq-B$ for some positive constants $A$ and $B$. Then Kobayashi metric and Carathéodory metric satisfy
@@
Here $C>0$ is depending only $A,B$, and independent of complex dimension $n$ of complex manifold $\mathcal{D}$.
@@
(cf. Theorem 4.2) For a suitable choice of positive constant $C$, suppose that $(M,h)$ is a complete Kähler manifold whose holomorphic sectional curvature $K_{h}$ satisfies $A\leq K_{h}\leq B$ for some positive constants $A$ and $B$. If the Carathéodory metric $\mathfrak{C}^{2}$ on complex manifold $M$ is a smooth strongly pseudoconvex complex Finsler metric. Then $H=C\mathfrak{C}^{2}+h$ and the Kähler metric $h$ are equivalent. What’s more, $M$ is a complete Kobayashi-hyperbolic manifold.
@@ 2 Terminology and preliminaries
In this section, we recall the following some basic definitions, facts and propositions in this paper. For more details, we refer to [1, 13, 5, 4].
2.1 Complex Finsler geometry
@@
is positive definite on $\widetilde{M}$.
@@
Any $C^{\infty}$ Hermitian metric on a complex manifold $M$ is naturally a strongly pseudoconvex complex Finsler metric. Conversely, if a complex Finsler metric $G$ on a complex manifold $M$ is $C^{\infty}$ over the whole holomorphic tangent bundle $T^{1,0}M$, then it is necessary a $C^{\infty}$ Hermitian metric. For this reason, in general the non-trivial (non-Hermitian quadratic) examples of complex Finsler metrics are only required to be smooth over the slit holomorphic tangent bundle $\widetilde{M}$.
If $\lambda=\lambda(z)dz\otimes d{\bar{z}},\lambda(z)>0$ is a Hermitian metric on a Riemann surface $R$. Then the Gaussian curvature $K(\lambda)(z)$ of $\lambda$ at the point $z$ is given by
$K(\lambda)(z)=-\frac{1}{\lambda(z)}\frac{\partial^{2}\log\lambda}{\partial z\partial\bar{z}}.$
Proposition 2.1.
For a positive constant $C>0$, let $\lambda$ be a Hermitian metric on a Riemann surface $R$. Then
$K(C\lambda)(z)=\frac{1}{C}K(\lambda)(z).$
Proof.
$\begin{split}K(C\lambda)(z)&=-\frac{1}{C\lambda(z)}\frac{\partial^{2}\log(C\lambda(z))}{\partial z\partial\bar{z}}\\ &=-\frac{1}{C\lambda(z)}\frac{\partial^{2}(\log\lambda(z)+\log C)}{\partial z\partial\bar{z}}\\ &=-\frac{1}{C}\frac{1}{\lambda(z)}\frac{\partial^{2}\log\lambda(z)}{\partial z\partial\bar{z}}\\ &=\frac{1}{C}K(\lambda)(z).\end{split}$
The following proposition was actually outlined and used in Wong and Wu [29, Lemma 4.3, p. 430]. We derive the proposition needed for Theorem 3.4 and give a brief proof here.
Proposition 2.2.
Suppose that $\lambda,\mu$ are Hermitian metrics on a Riemann surface $R$. Then
$K(\lambda+\mu)(z)\leq\frac{\lambda^{2}}{(\lambda+\mu)^{2}}K(\lambda)(z)+\frac{\mu^{2}}{(\lambda+\mu)^{2}}K(\mu)(z),$
where $K(\lambda)(z),K(\mu)(z),K(\lambda+\mu)(z)$ are the Gaussian curvature of the Hermitian metric $\lambda,\mu,\lambda+\mu$ at the point $z$, respectively. Moreover,
$K(\lambda+\mu)(z)\leq K(\lambda)(z)+K(\mu)(z).$
Proof.
By direct calculation, we have
$\begin{split}\frac{\partial^{2}\log(\lambda+\mu)}{\partial z\partial\bar{z}}&=\frac{\partial}{\partial z}\Big(\frac{\frac{\partial\lambda}{\partial\bar{z}}+\frac{\partial\mu}{\partial\bar{z}}}{\lambda+\mu}\Big)\\ &=\frac{\lambda\frac{\partial^{2}\lambda}{\partial z\partial\bar{z}}-\frac{\partial\lambda}{\partial z}\frac{\partial\lambda}{\partial\bar{z}}+\mu\frac{\partial^{2}\mu}{\partial z\partial\bar{z}}-\frac{\partial\mu}{\partial z}\frac{\partial\mu}{\partial\bar{z}}+\lambda\frac{\partial^{2}\mu}{\partial z\partial\bar{z}}+\mu\frac{\partial^{2}\lambda}{\partial z\partial\bar{z}}-\frac{\partial\lambda}{\partial z}\frac{\partial\mu}{\partial\bar{z}}-\frac{\partial\mu}{\partial z}\frac{\partial\lambda}{\partial\bar{z}}}{(\lambda+\mu)^{2}}\\ &=\frac{\lambda^{2}\frac{\partial^{2}\log\lambda}{\partial z\partial\bar{z}}+\mu^{2}\frac{\partial^{2}\log\mu}{\partial z\partial\bar{z}}+\frac{\lambda}{\mu}(\mu^{2}\frac{\partial^{2}\log\mu}{\partial z\partial\bar{z}}+\frac{\partial\mu}{\partial z}\frac{\partial\mu}{\partial\bar{z}})}{(\lambda+\mu)^{2}}\\ &+\frac{\frac{\mu}{\lambda}(\lambda^{2}\frac{\partial^{2}\log\lambda}{\partial z\partial\bar{z}}+\frac{\partial\lambda}{\partial z}\frac{\partial\lambda}{\partial\bar{z}})-\frac{\partial\lambda}{\partial z}\frac{\partial\mu}{\partial\bar{z}}-\frac{\partial\mu}{\partial z}\frac{\partial\lambda}{\partial\bar{z}}}{(\lambda+\mu)^{2}}\\ &=\frac{\lambda^{2}\frac{\partial^{2}\log\lambda}{\partial z\partial\bar{z}}+\mu^{2}\frac{\partial^{2}\log\mu}{\partial z\partial\bar{z}}+\lambda\mu(\frac{\partial^{2}\log\mu}{\partial z\partial\bar{z}}+\frac{\partial\log\mu}{\partial z}\frac{\partial\log\mu}{\partial\bar{z}})}{(\lambda+\mu)^{2}}\\ &+\frac{\lambda\mu(\frac{\partial^{2}\log\lambda}{\partial z\partial\bar{z}}+\frac{\partial\log\lambda}{\partial z}\frac{\partial\log\lambda}{\partial\bar{z}})-\lambda\mu\frac{\partial\log\lambda}{\partial z}\frac{\partial\log\mu}{\partial\bar{z}}-\lambda\mu\frac{\partial\log\mu}{\partial z}\frac{\partial\log\lambda}{\partial\bar{z}}}{(\lambda+\mu)^{2}}\\ &=\frac{\lambda^{2}\frac{\partial^{2}\log\lambda}{\partial z\partial\bar{z}}+\mu^{2}\frac{\partial^{2}\log\mu}{\partial z\partial\bar{z}}+\lambda\mu\frac{\partial^{2}\log\mu}{\partial z\partial\bar{z}}+\lambda\mu\frac{\partial^{2}\log\lambda}{\partial z\partial\bar{z}}}{(\lambda+\mu)^{2}}\\ &+\frac{\lambda\mu(\frac{\partial\log\lambda}{\partial z}-\frac{\partial\log\mu}{\partial z})(\frac{\partial\log\lambda}{\partial\bar{z}}-\frac{\partial\log\mu}{\partial\bar{z}})}{(\lambda+\mu)^{2}}\\ &=\frac{\lambda(\lambda+\mu)\frac{\partial^{2}\log\lambda}{\partial z\partial\bar{z}}+\mu(\lambda+\mu)\frac{\partial^{2}\log\mu}{\partial z\partial\bar{z}}+\lambda\mu|\frac{\partial\log\lambda}{\partial z}-\frac{\partial\log\mu}{\partial z}|^{2}}{(\lambda+\mu)^{2}}\end{split}$
Therefore we have
$\begin{split}&-\frac{\partial^{2}\log(\lambda+\mu)}{\partial z\partial\bar{z}}\leq-\frac{\lambda}{\lambda+\mu}\frac{\partial^{2}\log\lambda}{\partial z\partial\bar{z}}-\frac{\mu}{\lambda+\mu}\frac{\partial^{2}\log\mu}{\partial z\partial\bar{z}},\end{split}$
By the definition of Gaussian curvature, we obtain
$K(\lambda+\mu)(z)\leq\frac{\lambda^{2}}{(\lambda+\mu)^{2}}K(\lambda)(z)+\frac{\mu^{2}}{(\lambda+\mu)^{2}}K(\mu)(z)\leq K(\lambda)(z)+K(\mu)(z).$
@@
Let $(M,G)$ be a strongly pseudoconvex complex Finsler manifold $M$, and take $v\in\widetilde{M}$. Then the holomorphic sectional curvature $K_{G}(v)$ of $G$ along a non zero tangent vector $v$ is given by
@@
Abate and Patrizio found an phenomenon that the holomorphic sectional curvature of a complex Finsler metric $G$ is the supremum of the Gaussian curvature of the induced metric through a family of holomorphic maps.
@@
[…]
2.2 Invariant metrics
The Bergman pseudometric $g_{B}$ on a complex manifold $M$ of complex dimension $n$ is a Kähler pseudometric with local potential given by the coefficients of the Bergman kernel $K(x,x)$. Let $f$ be a $L^{2}$-holomorphic $n$-form on $M$. In terms of local coordinates $\left(z_{1},\ldots,z_{n}\right)$ on a coordinate chart $U$, let $e_{K_{M}}=dz^{1}\wedge\ldots\wedge dz^{n}$ be a local basis of the canonical line bundle $K_{M}$ on $U$. We can write $f$ as $f_{U}e_{K_{M}}$ on $U$. Let $f_{i}$ be an orthonormal basis of $L^{2}$-sections in $H_{(2)}^{0}\left(M,K_{M}\right)$. Note that from conformality, the choice is independent of the metric on $M$. The Bergman kernel is given by $K(x,x)=\sum_{i}f_{i}\wedge\overline{f_{i}}$. Let $K_{U}(x,x)=\sum_{i}f_{U,i}\overline{f_{U,i}}$ be the coefficient of $K(x,x)$ in terms of the local coordinates. The Bergman metric is given by the Kähler form